Consider the reaction between 60.0 ml of liquid methyl alcohol, CH3OH (density = .850 g/ml) and 30.1 liters of O2 at 27°C and a pressure of 1.5 atm. The products of the reaction are CO2(g) and H2O(g). Calculate the number of grams of CO2 formed if the reaction goes to completion. The balanced equation is: 2CH3OH + 3O2 ------> 2CO2 + 4H2O


1
Expert's answer
2021-04-25T23:10:55-0400

Mass of methyl alcohol = 60 × 0.850 = 51g

Mole= mass/molar mass = 51/32 = 1.59mole


Mole of O2 = PV/RT = 1.5 ×30.1 /(0.082 × 300)

= 1.84 mol

By the balanced equation, the limiting reagent is O2

3O2 = 2CO2 = 2 × 44g CO2

3 mole = 88g CO2

1.84 mole = 88 × 1.84 / 3

= 53.97g CO2